OpenStudy (anonymous):

On a Wisconsin road map, Kirstin measured a distance of 11 1/8 (11.125) inches from Madison to Wausau. The scale of this map is 1 inch for every 13 miles. What is the actual distance between Madison to Wausau to the nearest mile?

OpenStudy (albert0898):

11 Whole Inches = 11 * 13 = 143 Miles .125 Inches = .125 * 13 = 1.625 Miles Combine the two numbers together. 143 Miles + 1.625 Miles = 144.625 Round to the nearest mile and you'll get 145 Miles!

OpenStudy (anonymous):

thank you !

OpenStudy (albert0898):

Anytime!